OpenStudy (anonymous):

Here's another probability problem. Elsie draws two cards at random without replacement from a set of cards numbered 1 to 20. what is the probability that she first selects the card numbered 4 and then an odd-numbered card greater than 7? A) 3/200 B) 3/190 C) 1/40 D) 1/20

OpenStudy (anonymous):

\[P(x_1) = 1/20\] only 1 card numbered 4, 20 cards left. \[P(x_2) = 6/19\] 6 odd number that are greater than 7, 19 cards left \[P(x_1) * P(x_2) = 1/20 * 6/19 = 1/10 * 3/19 = 3/190\]